A Barium nucleus with a charge 56 and a Krypton nucleus with a charge 36 fly apart from rest at an initial separation of 8 fm (8 x 10-15m). How much kinetic energy do the two nuclei share when they are far apart?

A) 900 eV

B) 1.8 MeV

C)235 KeV

D)360 MeV

E)4.5x10^20 eV
Answer #1

Concept - the initial potential energy converts to kinetic energy when they are far apart. So, the kinetic energy is equal to initial potential energy,

***********************************************************************************************
Check the answer and let me know immediately if you find something wrong or missing... I will rectify the mistakes asap if any
Answer #2

Calculation:

1. Key Concepts:

  • Initial potential energy (U_initial) between the two nuclei is converted entirely into kinetic energy (K_final) when they are far apart (since they start from rest).

  • Coulomb potential energy formula:

    U=kq1q2r

    Where:

    • (Coulomb's constant),

    • q1=56eq2=36e (charges of Barium and Krypton nuclei),

    • e=1.6×1019C (elementary charge),

    • r=8×1015m.

2. Calculate Potential Energy (U_initial):

U=(8.99×109)×(56×1.6×1019)×(36×1.6×1019)8×1015

Simplify:

U=8.99×109×56×36×(1.6×1019)28×1015U=8.99×56×36×2.56×10298×1015U=4.64×10258×1015=5.8×1011J

3. Convert Joules to Electron-Volts (eV):

1eV=1.6×1019J    K=5.8×10111.6×1019=3.625×108eVK=362.5MeV360MeV

 

Answer:

The correct answer is D) 360 MeV.


answered by: anonymous